# What is the nature of late infall streamers?

### v1

@@ 3.2 Accretion in a turbulent medium
The emergence of fall-back shock streamers in the previous section is the result of the encounter with an initially spherical gas cloudlet. Even though it has evolved to a more extended structure by the time it reaches the star, the initial condition is representative of a considerable over-density in the ISM with a symmetric shape. However, it remains an open question how frequent such encounters with clearly discernible cloudlets of gas are, and how well they describe the conditions of the large-scale environment of a star-forming region. Even if they were to serve as a reasonable approximation, we find that the resulting streamer is short-lived. If encounters of this type were the dominant cause for streamers, they should be much rarer than the observational sample suggests (Garufi et al. 2024; Villenave et al. 2024).
@@
While this idea can help to understand the possibility of fall-back streamers arising, the full turbulent picture is more complex. In the following, we discuss simulation results from multiple simulations with various resulting structures. We investigated how the arising structures change with $v_{\mathrm{sys}}$, $\sigma_{\mathrm{turb}}$, and $k_{\mathrm{max}}$. Additionally, we considered the influence of $\dot{M}_{\mathrm{sys}}$ in Appendix B. The numerical choices for these parameters are shown in Table 1.
@@
In our analysis, we find that the morphology of the emerging infall structures is predominantly dependent on the flux of material into the region where the ISM gas has low-enough velocity to be initially bound to the star, and the resulting interaction. Gas originating from farther out regions is too energetic to remain bound to the disk, so that no fall-back occurs and no fall-back streamers can emerge. Neglecting the spatial details of the turbulent velocity field, we approximate this region as a sphere with radius $R_{\mathrm{bound}}$, given by
@@ 3.2.1 High systemic velocity: Weakly bound gas accretion
The consequences are two-fold. First, the angular momentum flux through this small volume is the lowest of all simulations and variable on the smallest timescales (see the blue line in Fig. 4). Therefore, any fall-back of material does not produce strong arcs. Second, the average linear momentum flux direction also varies strongly and on short timescales. This means that the inflowing gas does not create distinct density enhancements at specific locations, but is smeared out across the volume, preventing the formation of a long-lived and bright BHL accretion tail, as well as streamers that are distinct from this tail.
@@ 3.2.2 Reduced systemic velocity: Multiple arced streamers
For fall-back streamers to arise naturally in a simulation with no initial density enhancement, it is essential that the gas encountering the disk remains bound to the star after the encounter, so that the angular momentum flux is visible as an arc. In simulation number 2, we reduce the systemic velocity by a factor of 2, $v_{\mathrm{sys}}=$0.5\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$, while keeping the ratio between turbulent velocity dispersion and systemic velocity constant, $\sigma_{\mathrm{turb}}/v_{\mathrm{sys}}=1$. As a result, we find $R_{\mathrm{bound}}=$1775\text{\,}\mathrm{AU}$\$.
It is expected that, analogously to the scenario in the previous section, significant variations in the infall direction occur, because the systemic velocity and turbulent velocity dispersion are equal. However, unlike previously, the variations are expected to occur over longer timescales due to the larger $R_{\mathrm{bound}}$. For the same reason, the angular momentum flux should be higher. Indeed, both expectations are fulfilled in the simulation. The angular momentum flux is up to two orders of magnitude higher than previously (see the purple line in Fig. 4). The direction of the linear momentum flux (see Fig. 5) and the angular momentum flux vary comparatively less strongly and on longer timescales.
@@
As the structures in this setup are highly time dependent due to the strong turbulence, we show a time series of the CO emission in four simulation snapshots in Fig. 8. At first, only weak non-Keplerian features arise around the disk, and the locations of the peak of the CO emission is not correlated with it. In the second panel, the first streamer on the left has formed, and the formation of the one on the right has started. Both big streamers are visible in the third panel, but a third one at the top also appears. It is distinct at this time, but merges with the right streamer by the time of the snapshot shown in Fig. 7. In the end, as seen in the fourth panel, the remaining structures are narrow, fainter, less extended inflows starting to dissipate. In a more realistic scenario with strong turbulence persisting for longer times, continued interactions may continuously lead to the creation of new, strong streamers as old ones dissipate. We therefore expect that fall-back streamers could arise abundantly in this type of environment, leading to an increased detection probability compared to the interaction with a single cloudlet.
@@ 3.2.3 Reduced turbulence: Single arced streamer
The emerging infall structures show two predominant features. One is the previously absent straight inflow originating from the BHL accretion tail, which can now be identified more clearly due to less severe turbulent fluctuations. It is supplied by the flow related the systemic velocity, which carries no angular momentum relative to the star-disk system. Consequently, the resulting feature is not arced and has a substantially extended morphology, so that the appearance of this accretion tail is less reminiscent of a typical streamer in scattered light and the CO moment maps (see Fig. 10). The intensity of the molecular emission is low, so that the accretion tail is very faint with only some areas with $T_{\mathrm{b,peak}}>$9\text{\,}\mathrm{K}$\$, and a distinction from the background in the peak brightness temperature map is difficult.
The second feature is an arc of gas, located at the bottom left edge of the disk. The moment 1 map of the CO emission reveals that it is an inflowing structure with a velocity that is different from the straight accretion tail, as shown in the middle panel of Fig. 10, confirming it as a fall-back streamer. Compared to the previously discussed case with higher turbulence, only a single, more strongly arced streamer arises in this scenario. Its spatial origin is located close to the straight accretion tail, which is expected due to the lower amplitude of the turbulent waves. The corresponding CO emission is dim, like the straight accretion tail, but the arc is encompassed by the $200\text{\,}\mathrm{mJy}\text{\,}{\mathrm{arcsec}}^{-2}\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$ moment 0 contour. During earlier times in this simulation, an additional arced streamer is visible for a short time interval, from ${\sim}$20\text{\,}\mathrm{kyr}$\$ to ${\sim}$25\text{\,}\mathrm{kyr}$\$. At this time, the qualitative inflow structure is similar to the bottom and left streamers found in simulation 2 (see Fig. 7), but the bottom streamer is spatially and dynamically close to the accretion tail as a result of the lower turbulence. As a result, it has dissipated at the time shown in Fig. 10.
@@ 4 Discussion
4.1 Fall-back streamers should be common
@@
Previous modeling of streamers that only consider them as originating from clear mass filaments, or as the signatures of the accretion of massive gas cloudlets (see Section 3.1), imply that streamers should be rare. This conclusion is strengthened further by the fact that streamer-like structures do not persist for a large fraction of the disk lifetime: only ${\sim}1\%$ of it assuming a conservative disk lifetime of $1\text{\,}\mathrm{Myr}$.
@@
Unless such encounters are very frequent to replenish the inflow and the related signatures, they fall short of explaining the high abundance of late-infall streamers suggested by observations (Gupta et al. 2023; Garufi et al. 2024). In this case, the fall-back streamers we find in the described simulations may offer an explanation. Here, streamers are a common phenomenon for disk-star systems moving through a dense environment. Even though individual streamers do not last considerably longer than in the case of a cloudlet encounter, their emergence does not depend on a specific encounter, but are the result of smaller asymmetries caused by turbulence. As a result, they should arise continuously, increasing the total time they may be observable. In fact, even a low level of turbulence can create a single arced streamer on top of the straight accretion tail.
@@
The fall-back streamers, even though their morphology is similar to a stream of material originating from a mass reservoir, do not hold the same information about the large-scale environment. In particular, the infall direction that could be obtained through dynamical modeling is not related to the spatial location of an over-density in the molecular cloud, and a direct connection may not be found. Even in the simplified case of cloudlet accretion, the fall-back streamers origin is different from the origin of the cloudlet initially encountering the disk. For turbulent accretion, especially in a highly turbulent environment, the orientation of the streamers is fully random. Here, only the straight infall directly from the accretion tail can be used to determine the systemic velocity of the disk.
@@
Infall of material via BHL accretion after the initial core collapse is also found in parsec-scale star-formation simulations (Padoan et al. 2025; Pelkonen et al. 2025). Previous studies attribute various protoplanetary disk features to this phenomenon (Hennebelle et al. 2017; Kuffmeier et al. 2018, 2023; Winter et al. 2024), further emphasizing its relevance and the need to more in-depth studies of its implications. In this work, we consider the occurrence and morphology of streamers arising in the BHL accretion scenario, and we explore a multitude of environmental conditions. Such considerations were generally not subject of previous large-scale simulations. We focus on the environment around a single star with a simplified treatment of the environment, identifying fall-back as the source for the streamers in our simulations. In the future, our findings should be compared to the morphology and occurrence of accretion structures around disks arising self-consistently in parsec-scale star-formation simulations. The majority of such simulations do not currently reach far into the Class II phase, though.
@@ 4.2 Streamer morphology as a constraint for local cloud conditions
The detection of a straight inflow signature, which can be identified as the accretion tail caused by the systemic flow, serves as strong evidence for a dense environment that may cause additional inflow signatures. The orientation of the straight accretion tail should then be correlated with the proper motion of the disk’s host star determined by, for example, Gaia. In contrast, the orientation of the fall-back streamers caused by turbulence should be uncorrelated. Detecting a single streamer would be indicative of a low turbulence environment, but if no corresponding BHL accretion tail is found, a formation scenario related to the cloudlet case may be the favored explanation, implying that the environmental density is low. It may then be caused by such a single encounter, or originate from a mass reservoir. We note that in environments with moderate levels of turbulence, like in our simulation number 4, the presence of an arced streamer may be missed by observations as emission is faint.
@@ 4.4 Model caveats
When computing the synthetic images, the freeze-out and photodissociation of CO was neglected. The freeze-out of CO onto interstellar dust is particularly relevant for the cold regions far from stellar heating sources. It could act to reduce the background emission, which would enhance the contrast of the streamer structures and therefore their observability. Conversely, the photodissociation of CO could reduce emission close to the star, balanced by shielding effects of molecular hydrogen and CO itself. To estimate the strength of this effect, we performed order-of-magnitude calculations of the CO column density, assuming photodissociation as the only reaction. We assume a moderately high unattenuated UV field strength $\chi=${10}^{5}$\$, given in units of the Draine field and at a reference radius of $100\text{\,}\mathrm{AU}$, used reaction rates from the UMIST database for astrochemisty (Woodall et al. 2007), and employed the self-shielding prescription from Woitke et al. (2009). We neglected extinction by the ISM dust. In regions with a lower density of ${10}^{-19}\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-3}$, 50% of CO at a distance of $1000\text{\,}\mathrm{AU}$ may be subject to dissociation over a streamer lifetime of $10\text{\,}\mathrm{kyr}$, but almost all CO inside $500\text{\,}\mathrm{AU}$ could be dissociated. In regions with a high density of ${10}^{-17}\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-3}$, only 2% of CO is dissociated close to the star at $200\text{\,}\mathrm{AU}$. Therefore, photodissociation may play a significant role for low-density streamers that arise in high-turbulence regions, but more detailed modeling would be required to accurately assess how astrochemical processes affect the observability of the streamers we find using CO lines.
@@ 5 Conclusions
  • The encounter of a Class II disk with an expanding cloudlet results in the formation of a ”fall-back” streamer, where material that initially overshot remains bound and falls back onto the disk with relative angular momentum.
@@
Our findings suggest that, while late infall with a clear origin, represented by cloudlet encounters, can create streamers, they can also arise naturally as an evolved protoplanetary disk moves through the turbulent ISM. This can reconcile ubiquitous detections of late-infall streamers with their short lifetime. Because their morphology is sensitive to the turbulent conditions of the environment and the accretion rate onto the disk, detailed studies of streamers may shed light on the conditions present in the environment of planet-forming disks, which may help to improve our understanding of planet formation.
@@ Acknowledgements.
We thank Jiahan Shi, Jorge Pérez González and Dmitry Semenov for insightful discussions. L.-A. H. acknowledges funding by the DFG via the Heidelberg Cluster of Excellence STRUCTURES in the framework of Germany’s Excellence Strategy (grant EXC-2181/1 – 390900948). We acknowledge support by the High Performance and Cloud Computing Group at the Zentrum für Datenverarbeitung of the University of Tübingen, the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grants INST 35/1134-1 FUGG, 35/1597-1 FUGG and 37/935-1 FUGG.
@@ Appendix A Turbulent velocity field
We model turbulence as a Gaussian random field, whose strength is defined by $\sigma_{\mathrm{turb}}$. Figure 12 shows the realization of this field for one velocity component, Cartesian $v_{z}$, for simulations number 1 and 5 from Table 1, which use $\sigma_{\mathrm{turb}}=$1\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$ and $\sigma_{\mathrm{turb}}=$0.1\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$, respectively. The resulting turbulent velocity fluctuations are less significant for simulation number 5, but not negligible. We note that the depicted velocity component contains the contribution from the systemic velocity of the disk.

### v2

@@ 3.2 Accretion in a turbulent medium
The emergence of fallback shock streamers in the previous section is the result of the encounter with an initially spherical gas cloudlet. Even though it has evolved to a more extended structure by the time it reaches the star, the initial condition is representative of a considerable over-density in the ISM with a symmetric shape. However, it remains an open question how frequent such encounters with clearly discernible cloudlets of gas are, and how well they describe the conditions of the large-scale environment of a star-forming region. Even if they were to serve as a reasonable approximation, we find that the resulting streamer is short-lived. If encounters of this type were the dominant cause for streamers, they should be much rarer than the observational sample suggests (Garufi et al. 2024; Villenave et al. 2024).
@@
While this idea can help to understand the possibility of fallback streamers arising, the full turbulent picture is more complex. In the following, we discuss simulation results from multiple simulations with various resulting structures. We investigated how the arising structures change with $v_{\mathrm{sys}}$, $\sigma_{\mathrm{turb}}$, and $k_{\mathrm{max}}$. We considered the influence of $\dot{M}_{\mathrm{sys}}$ in Appendix B. The numerical choices for these parameters are shown in Table 1.
@@
In our analysis, we find that the morphology of the emerging infall structures is predominantly dependent on the flux of material into the region where the ISM gas has low-enough velocity to be initially bound to the star, and the resulting interaction. Gas originating from farther out regions is too energetic to remain bound to the disk, so that no fallback occurs and no fallback streamers can emerge. Neglecting the spatial details of the turbulent velocity field, we approximated this region as a sphere with radius $R_{\mathrm{bound}}$, given by
@@ 3.2.1 High systemic velocity: Weakly bound gas accretion
The consequences are two-fold. First, the angular momentum flux through this small volume is the lowest of all simulations and variable on the smallest timescales (see the blue line in Fig. 4). Therefore, any fallback of material does not produce strong arcs. Second, the average linear momentum flux direction also varies strongly and on short timescales. This means that the inflowing gas does not create distinct density enhancements at specific locations, but is smeared out across the volume, preventing the formation of a long-lived and bright BHL accretion tail, as well as streamers that are distinct from this tail.
@@ 3.2.2 Reduced systemic velocity: Multiple arced streamers
For fallback streamers to arise naturally in a simulation with no initial density enhancement, it is essential that the gas encountering the disk remains bound to the star after the encounter, so that the angular momentum flux is visible as an arc. In simulation number 2, we reduced the systemic velocity by a factor of 2, $v_{\mathrm{sys}}=$0.5\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$, while keeping the ratio between turbulent velocity dispersion and systemic velocity constant, $\sigma_{\mathrm{turb}}/v_{\mathrm{sys}}=1$. As a result, we find $R_{\mathrm{bound}}=$1775\text{\,}\mathrm{AU}$\$.
We expect that, analogously to the scenario in the previous section, significant variations in the infall direction occur, because the systemic velocity and turbulent velocity dispersion are equal. However, unlike previously, the variations should occur over longer timescales due to the larger $R_{\mathrm{bound}}$. For the same reason, the angular momentum flux should be higher. Indeed, both expectations are fulfilled in the simulation. The angular momentum flux is up to two orders of magnitude higher than previously (see the purple line in Fig. 4). The direction of the linear momentum flux (see Fig. 5) and the angular momentum flux vary comparatively less strongly and on longer timescales.
@@
As the structures in this setup are highly time dependent due to the strong turbulence, we show a time series of the CO emission in four simulation snapshots in Fig. 8. At first, only weak non-Keplerian features arise around the disk, and the locations of the peak of the CO emission is not correlated with it. In the second panel, the first streamer on the left has formed, and the formation of the one on the right has started. Both big streamers are visible in the third panel, but a third one at the top also appears. It is distinct at this time, but merges with the right streamer by the time of the snapshot shown in Fig. 7. In the end, as seen in the fourth panel, the remaining structures are narrow, fainter, less extended inflows starting to dissipate. In a more realistic scenario with strong turbulence persisting for longer times, continued interactions may continuously lead to the creation of new, strong streamers as old ones dissipate. We therefore expect that fallback streamers could arise abundantly in this type of environment, leading to an increased detection probability compared to the interaction with a single cloudlet.
@@ 3.2.3 Reduced turbulence: Single arced streamer
The emerging infall structures show two predominant features. One is the previously absent straight inflow originating from the BHL accretion tail, which can now be identified more clearly due to less severe turbulent fluctuations. It is supplied by the flow related to the systemic velocity, which carries no angular momentum relative to the star-disk system. Consequently, the resulting feature is not arced and has a substantially extended morphology, so that the appearance of this accretion tail is less reminiscent of a typical streamer in scattered light and the CO moment maps (see Fig. 10). The intensity of the molecular emission is low, so that the accretion tail is very faint with only some areas with $T_{\mathrm{b,peak}}>$9\text{\,}\mathrm{K}$\$, and a distinction from the background in the peak brightness temperature map is difficult.
The second feature is an arc of gas, located at the bottom left edge of the disk. The moment 1 map of the CO emission reveals that it is an inflowing structure with a velocity that is different from the straight accretion tail, as shown in the middle panel of Fig. 10, confirming it as a fallback streamer. Compared to the previously discussed case with higher turbulence, only a single, more strongly arced streamer arises in this scenario. Its spatial origin is located close to the straight accretion tail, which is expected due to the lower amplitude of the turbulent waves. The corresponding CO emission is dim, like the straight accretion tail, but the arc is encompassed by the $200\text{\,}\mathrm{mJy}\text{\,}{\mathrm{arcsec}}^{-2}\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$ moment 0 contour. During earlier times in this simulation, an additional arced streamer is visible for a short time interval, from ${\sim}$20\text{\,}\mathrm{kyr}$\$ to ${\sim}$25\text{\,}\mathrm{kyr}$\$. At this time, the qualitative inflow structure is similar to the bottom and left streamers found in simulation 2 (see Fig. 7), but the bottom streamer is spatially and dynamically close to the accretion tail as a result of the lower turbulence. As a result, it has dissipated at the time shown in Fig. 10.
@@ 4 Discussion
4.1 Fallback streamers should be common
@@
Previous modeling of streamers that only consider them as originating from clear mass filaments, or as the signatures of the accretion of massive gas cloudlets (see Section 3.1), implies that streamers should be rare. This conclusion is strengthened further by the fact that streamer-like structures do not persist for a large fraction of the disk lifetime: only ${\sim}1\%$ of it assuming a conservative disk lifetime of $1\text{\,}\mathrm{Myr}$.
@@
Unless such encounters are very frequent to replenish the inflow and the related signatures, they fall short of explaining the high abundance of late-infall streamers suggested by observations (Gupta et al. 2023; Garufi et al. 2024). In this case, the fallback streamers we find in the described simulations may offer an explanation. Here, streamers are a common phenomenon for disk-star systems moving through a dense environment. Even though individual streamers do not last considerably longer than in the case of a cloudlet encounter, their emergence does not depend on a specific encounter, but are the result of smaller asymmetries caused by turbulence. As a result, they should arise continuously, increasing the total time they may be observable. In fact, even a low level of turbulence can create a single arced streamer on top of the straight accretion tail.
@@
The fallback streamers, even though their morphology is similar to a stream of material originating from a mass reservoir, do not hold the same information about the large-scale environment. In particular, the infall direction that could be obtained through dynamical modeling is not related to the spatial location of an over-density in the molecular cloud, and a direct connection may not be found. Even in the simplified case of cloudlet accretion, the fallback streamers origin is different from the origin of the cloudlet initially encountering the disk. For turbulent accretion, especially in a highly turbulent environment, the orientation of the streamers is fully random. Here, only the straight infall directly from the accretion tail can be used to determine the systemic velocity of the disk.
@@
Infall of material via BHL accretion after the initial core collapse is also found in parsec-scale star-formation simulations (Padoan et al. 2025; Pelkonen et al. 2025). Previous studies attribute various protoplanetary disk features to this phenomenon (Hennebelle et al. 2017; Kuffmeier et al. 2018, 2023; Winter et al. 2024), further emphasizing its relevance and the need to more in-depth studies of its implications. In this work, we considered the occurrence and morphology of streamers arising in the BHL accretion scenario, and we explored a multitude of environmental conditions. Such considerations were generally not subject of previous large-scale simulations. We focussed on the environment around a single star with a simplified treatment of the environment, identifying fallback as the source for the streamers in our simulations. In the future, our findings should be compared to the morphology and occurrence of accretion structures around disks arising self-consistently in parsec-scale star-formation simulations. The majority of such simulations do not currently reach far into the Class II phase, though.
@@ 4.2 Streamer morphology as a constraint for local cloud conditions
The detection of a straight inflow signature, which can be identified as the accretion tail caused by the systemic flow, serves as strong evidence for a dense environment that may cause additional inflow signatures. The orientation of the straight accretion tail should then be correlated with the proper motion of the disk’s host star determined by, for example, Gaia. In contrast, the orientation of the fallback streamers caused by turbulence should be uncorrelated. Detecting a single streamer would be indicative of a low turbulence environment, but if no corresponding BHL accretion tail is found, a formation scenario related to the cloudlet case may be the favored explanation, implying that the environmental density is low. It may then be caused by such a single encounter, or originate from a mass reservoir. We note that in environments with moderate levels of turbulence, like in our simulation number 4, the presence of an arced streamer may be missed by observations as emission is faint.
@@ 4.4 Model caveats
When computing the synthetic images, the freeze-out and photodissociation of CO was neglected. The freeze-out of CO onto interstellar dust is particularly relevant for the cold regions far from stellar heating sources. It could reduce the background emission, which would enhance the contrast of the streamer structures and thereby their observability. Conversely, the photodissociation of CO could reduce emission close to the star, balanced by shielding effects of molecular hydrogen and CO itself. To estimate the strength of this effect, we performed order-of-magnitude calculations of the CO column density, assuming photodissociation as the only reaction. We assumed a moderately high unattenuated UV field strength $\chi=${10}^{5}$\$, given in units of the Draine field and at a reference radius of $100\text{\,}\mathrm{AU}$, used reaction rates from the UMIST database for astrochemistry (Woodall et al. 2007), and employed the self-shielding prescription from Woitke et al. (2009). We neglected extinction by the ISM dust. In regions with a lower density of ${10}^{-19}\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-3}$, 50% of CO at a distance of $1000\text{\,}\mathrm{AU}$ may dissociate over a streamer lifetime of $10\text{\,}\mathrm{kyr}$, but almost all CO inside $500\text{\,}\mathrm{AU}$ could be dissociated. In regions with a high density of ${10}^{-17}\text{\,}\mathrm{g}\text{\,}{\mathrm{cm}}^{-3}$, only 2% of CO is dissociated close to the star at $200\text{\,}\mathrm{AU}$. Therefore, photodissociation may play a significant role for low-density streamers arising in high-turbulence regions, but more detailed modeling would be required to accurately assess how astrochemical processes affect the observability of the streamers we find using CO lines.
@@ 5 Conclusions
  • The encounter of a Class II disk with an expanding cloudlet results in the formation of a ”fallback” streamer, where material that initially overshot remains bound and falls back onto the disk with relative angular momentum.
@@
Our findings suggest that while late infall with a clear origin, represented by cloudlet encounters, can create streamers, they can also arise naturally as an evolved protoplanetary disk moves through the turbulent ISM. This can reconcile ubiquitous detections of late-infall streamers with their short lifetime. Because their morphology is sensitive to the turbulent conditions of the environment and the accretion rate onto the disk, detailed studies of streamers may shed light on the conditions present in the environment of planet-forming disks, which may help to improve our understanding of planet formation.
@@ Acknowledgements.
We thank Jiahan Shi, Jorge Pérez González, and Dmitry Semenov for insightful discussions. L.-A. H. acknowledges funding by the DFG via the Heidelberg Cluster of Excellence STRUCTURES in the framework of Germany’s Excellence Strategy (grant EXC-2181/1 – 390900948). We acknowledge support by the High Performance and Cloud Computing Group at the Zentrum für Datenverarbeitung of the University of Tübingen, the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grants INST 35/1134-1 FUGG, 35/1597-1 FUGG and 37/935-1 FUGG.
@@ Appendix A Turbulent velocity field
We modeled turbulence as a Gaussian random field, whose strength is defined by $\sigma_{\mathrm{turb}}$. Figure 12 shows the realization of this field for one velocity component, Cartesian $v_{z}$, for simulations number 1 and 5 from Table 1, which use $\sigma_{\mathrm{turb}}=$1\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$ and $\sigma_{\mathrm{turb}}=$0.1\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$, respectively. The resulting turbulent velocity fluctuations are less significant for simulation number 5, but not negligible. We note that the depicted velocity component contains the contribution from the systemic velocity of the disk.